can you break down this word problem for me: You need to purchase a textbook for nursing school. The book cost $80.00, and the sales tax where you are purchasing the book is 8.25%. You have $100. How much change will you receive back?
Answer by stanbon(75887)

can you break down this word problem for me: You need to purchase a textbook for nursing school. The book cost $80.00, and the sales tax where you are purchasing the book is 8.25%. You have $100. How much change will you receive back?
----
Change = 100 -(80 + 0.0825*80)
----
Change = 100 - (1.0825*80)
----
Change = 100 - 86.60
----
change = $13.40
==========================
Cheers,
Stan H.